# Efficient Large Language Models: A Survey

### v2

@@ 2.5.4. Transformer-Alternate Architectures
State Space Models. A promising approach that aims to substitute the attention mechanism is state space models (SSMs). SSM is formulated as , , which maps a single-dimension input signal to an N-dimension latent state before projecting to a single-dimension output signal , where , , , are parameters learned by gradient descent (Gu et al., 2022a). Compared to attention that has quadratic complexity, SSMs provide near-linear computational complexity relative to the length of the sequence. Given such advantage, a series of techniques have been proposed to improve SSMs. For example, the Structured State Space sequence model (S4) (Gu et al., 2022a) refines SSMs by conditioning matrix with a low-rank correction. This enables stable diagonalization and simplifies the SSM to the well-studied computation of a Cauchy kernel. Diagonal State Space (DSS) (Gupta et al., 2022) improves SSMs by proposing fully diagonal parameterization of state spaces instead of a diagonal plus low rank structure, demonstrating greater efficiency. To bridge the gap between SSMs and attention while adapting to modern hardware, H3 (Fu et al., 2023b) stacks two SSMs to interact with their output and input projection, allowing it to log tokens and facilitate sequence-wide comparisons simultaneously. Mehta et al. (2022) introduce a more efficient layer called Gated State Space (GSS), which has been empirically shown to be 2-3 times faster than the previous strategy (Gupta et al., 2022) while maintaining the perplexity on multiple language modeling benchmarks. Block-State Transformer (BST) (Pilault et al., 2023) designs a hybrid layer that combines an SSM sublayer for extended range contextualization with a Block Transformer sublayer for short-term sequence representation. Gu and Dao (2023) propose Mamba to enhance SSMs by designing a selection mechanism to eliminate irrelevant data and developed a hardware-aware parallel algorithm for recurrent operation, achieving 5x higher throughput than Transformers. Ren et al. (2023a) propose a general modular activation mechanism, Sparse Modular Activation (SMA), that unifies previous works on MoE, adaptive computation, dynamic routing and sparse attention, and further applies SMA to develop a novel architecture, SeqBoat, to achieve state-of-the-art quality-efficiency trade-off.
@@
Other Sequential Models. Lastly, some other architectures have been proposed to replace the Transformer layer. Receptance Weighted Key Value (RWKV) model (Peng et al., 2023a) amalgamates the advantages of recurring neural networks (RNN) and Transformers. This combination is designed to utilize the effective parallelizable training feature of Transformers coupled with the efficient inference ability of RNNs, thereby forging a model adept at managing auto-regressive text generation and effectively tackling challenges associated with long sequence processing. Poli et al. (2023) propose Hyena, a sub-quadratic alternative to the attention mechanism, mitigating the quadratic cost in long sequences. This operator includes two efficient sub-quadratic primitives: an implicit long convolution and multiplicative element-wise gating of the input. Through this, Hyena facilitates the development of larger, more efficient convolutional language models for long sequences. Lastly, MEGABYTE (Yu et al., 2023) breaks down long byte sequences into fixed-sized patches akin to tokens, comprising a patch embedder for encoding, a global module acting as a large autoregressive Transformer for patch representations, and a local module for predicting bytes within a patch.
3. Data-Centric Methods
@@
where level=1text width=19em,font=,,
@@
Yao
et al. (2022b),
DSIR (Xie
et al., 2023), leaf, text width=17em
@@
Instruction Mining (Cao
et al., 2023),
Ivison
et al. (2023),
TS-DShapley (Schoch
et al., 2023),
LTD Instruction Tuning (Chen et al., 2023i),
AlpaGasus (Chen et al., 2023e),
LIMA (Zhou et al., 2023a), leaf, text width=27em
3.1. Data Selection
Data selection for LLMs involves carefully selecting the most informative and diverse examples so that the model can efficiently capture essential patterns and features, accelerating the learning process (Xie et al., 2023; Yao et al., 2022b; Santamaría and Axelrod, 2019; Glass et al., 2020). Figure 15 summarizes the latest data selection techniques for efficient LLM pre-training and fine-tuning.
3.1.1. Data Selection for Efficient Pre-Training
@@
Data selection enhances LLMs pre-training efficiency by allowing the model to focus on the most informative and relevant examples during training. By carefully curating a subset of representative data, the model can extract essential patterns and features, leading to a more efficient acquisition of generalized knowledge. For example, SSPT (Glass et al., 2020) is a pre-training task based on the principles of reading comprehension. It involves selecting answers from contextually relevant text passages, which has shown notable improvements in performance across various Machine Reading Comprehension (MRC) benchmarks. Yao et al. (2022b) propose a meta-learning-based method for the selection of linguistically informative sentences which significantly elevates the quality of machine-generated translations. Xie et al. (2023) propose DSIR, a data selection method based on importance re-sampling for both general-purpose and specialized LLMs. It calculates how important different pieces of data are within a simpler set of features and chooses data based on these importance calculations.
3.1.2. Data Selection for Efficient Fine-Tuning
@@
Data selection can also boost fine-tuning efficiency since only a curated subset of examples is employed to refine the model. This approach ensures that the adaptation process is conducted with a focus on the specific nuances intrinsic to the target domain or task, making the fine-tuning process more efficient. For example, Instruction Mining (Cao et al., 2023) presents a linear evaluation method to assess data quality in instruction-following tasks. It highlights the importance of high-quality data, showing that models trained with Instruction Mining-curated datasets outperform those trained on generic datasets in 42.5% of cases. This underscores the significance of data quality and lays the groundwork for future improvements in instruction-following model efficacy. Ivison et al. (2023) propose to use a few unlabeled examples to retrieve similar labeled ones from a larger multitask dataset, improving task-specific model training. This method outperforms standard multitask data sampling for fine-tuning and enhances few-shot fine-tuning, yielding a 2-23% relative improvement over current models. TS-DShapley (Schoch et al., 2023) is introduced to address the computational challenges of applying Shapley-based data valuation to fine-tuning LLMs. It employs an efficient sampling-based method that aggregates Shapley values computed from subsets to evaluate the entire training set. Moreover, it incorporates a value transfer method that leverages information from a simple classifier trained using representations from the target language model. Low Training Data Instruction Tuning (LTD Instruction Tuning) (Chen et al., 2023i) challenges the need for large datasets in fine-tuning, showing that less than 0.5% of the original dataset can effectively train task-specific models without compromising performance. This approach enables more resource-efficient practices in data-scarce environments, combining selective data strategies with tailored training protocols for optimal data efficiency. AlpaGasus (Chen et al., 2023e) is a model fine-tuned on a mere 9k high-quality data points, which are meticulously filtered from a larger dataset of 52k. It outperforms the original model trained on the full dataset and reduces training time by 5.7x, demonstrating the power of high-quality data in instruction-fine-tuning. LIMA (Zhou et al., 2023a) fine-tunes LLMs with a small, selected set of examples, showing strong performance and challenging the need for extensive tuning. It generalizes well to new tasks and, in comparisons, matches or exceedes GPT-4 in 43% of cases, suggesting that LLMs gain most knowledge in pre-training, requiring minimal instruction tuning.
@@
where level=1text width=10em,font=,,
where level=2text width=13.5em,font=,,
where level=3text width=12em,font=,,
@@
KATE (Liu
et al., 2022a),
VoteK (Su et al., 2022a),
Wang
et al. (2023i),
IDS (Qin
et al., 2023b),
Min et al. (2022),
LENS (Li and Qiu, 2023),
MDL (Wu
et al., 2023b),
Zhang
et al. (2022b)
EPR (Rubin
et al., 2022),
UDR (Li et al., 2023c),
Wang
et al. (2023g),
@@
, leaf, text width=24em
@@
Liu
et al. (2022a),
Lu
et al. (2022a),leaf, text width=13em
@@
Instruction Induction (Honovich
et al., 2022),
Automatic Prompt Engineer (Zhou et al., 2023b),
Self-Instruct (Wang et al., 2023b),
OPRO (Yang
et al., 2023c),
TeGit (Chen
et al., 2023c)
, leaf, text width=28em
@@
ReAct (Yao et al., 2023e),
Least-to-Most Prompting (Zhou et al., 2023c),
Tree-of-Thought (Yao et al., 2023c),
CoT-SC (Wang et al., 2022e),
Graph of Thoughts (Besta et al., 2023),
Contrastive CoT (Chia
et al., 2023),
XoT (Ding et al., 2023b), leaf, text width=26em
@@
Gisting (Mu et al., 2023), AutoCompressors (Chevalier
et al., 2023),
PCRL (Jung and Kim, 2023),
ICAE (Ge
et al., 2023),
Nugget 2D (Qin
et al., 2023a),
LongLLMLingua (Jiang
et al., 2023c), leaf, text width=43em
@@
TempLM (Zhang
et al., 2022c),
PromptGen (Zhang
et al., 2022a),
leaf, text width=25em
3.2. Prompt Engineering
Prompt engineering (Liu et al., 2023g) focuses on designing effective inputs (i.e., prompts) to guide LLMs in generating desired outputs. It enhances inference efficiency by tailoring the input prompts or queries to better suit the capabilities and nuances of a specific language model. When used for some simple tasks, such as semantic classification, prompt engineering can even substitute fine-tuning to achieve high accuracy (Liu et al., 2022c). As summarized in Figure 17, prompt engineering techniques can be grouped into few-shot prompting, prompt compression, and prompt generation.
3.2.1. Few-Shot Prompting
Few-shot prompting involves providing a LLM with a limited set of examples (i.e., demonstrations) to steer its understanding to a task it is required to execute (Wei et al., 2022a). These demonstrations are selected from the LLM’s training corpus based on their similarity to the test example, and the LLM is expected to use the knowledge gained from these similar demonstrations to make the correct prediction (Dong et al., 2023). Few-shot prompting provides an efficient mechanism to use LLM by guiding the LLM to perform a wide variety of tasks without the need for additional training or fine-tuning. Furthermore, an effective few-shot prompting approach can make the created prompt concise enough to allow LLMs to quickly adjust to the task in high accuracy with only a slight increase of extra context, thus significantly improving inference speed. As illustrated in Figure 18, few-shot prompting techniques can generally be grouped into demonstration selection, demonstration ordering, instruction generation, and multi-step reasoning.
@@ 3.2.1. Few-Shot Prompting
Demonstration Organization. Demonstration organization refers to organizing the demonstrations in an appropriate way so as to form a suitable prompt for inference. Demonstration organization has a significant impact on the inference speed. Improper organization may result in the processing of a considerable amount of unnecessary information, leading to significant slowdown. The main challenges of demonstration organization come from two perspectives: demonstration selection and demonstration ordering.
@@
Demonstration Selection. Demonstration selection aims to choose the good examples for few-shot prompting (Dong et al., 2023). In order to generate a satisfactory result, a good selection of demonstrations may only require a few number of demonstrations to be used for the prompt, thus making the prompt concise and straightforward for a more efficient inference. Existing demonstration selection techniques can be grouped into unsupervised methods (Liu et al., 2022a; Su et al., 2022a; Wang et al., 2023i; Qin et al., 2023b; Min et al., 2022; Li and Qiu, 2023; Wu et al., 2023b; Zhang et al., 2022b) and supervised methods (Rubin et al., 2022; Li et al., 2023c; Wang et al., 2023g; Luo et al., 2023). Unsupervised methods aim to select the nearest examples from the training set using a predefined similarity function, such as L2 distance, cosine distance, and the minimum description length (MDL) (Wu et al., 2023b). For example, KATE (Liu et al., 2022a) is an unsupervised selection method that directly uses the nearest neighbors of a given test sample as the corresponding demonstrations. VoteK (Su et al., 2022a) is an improved version of KATE to resolve its limitation that requires a large set of examples to achieve good performance. Unlike KATE, VoteK increases the diversity of the demonstrations by penalizing examples similar to those already selected. In comparison, supervised methods require training a domain-specific retriever from the training set and using it for demonstration selection. For example, EPR (Rubin et al., 2022) is trained to select demonstrations from a small set of candidates initialized by the unsupervised retriever such as BM25 from the training corpse. UDR (Li et al., 2023c) further enhances EPR by adopting a unified demonstration retriever to unify the demonstration selection across different tasks. Compared to unsupervised methods, supervised methods often lead to a more satisfying generation result but require frequent adjustment of the retriever for handling the out-of-domain data, making them less efficient for inference.
@@
Demonstration Ordering. After selecting representative samples from the training set, the next step is ordering these samples in the prompt. The order of the demonstrations also has a significant impact on the performance of the model. Therefore, selecting the right order of demonstrations can help the model quickly reach a good generation quality with fewer samples, thus improving the inference efficiency. To date, only a few studies have delved into this area. For example, Liu et al. (2022a) suggest arranging demonstrations based on their distance from the input, placing the closest demonstration furthest to the right. Lu et al. (2022a) propose to develop both global and local entropy metrics and use the entropy metrics to set up the demonstration order.
@@
Template Formatting. Template Formatting aims to design a suitable template to form the prompt. A good template typically compiles all the information needed by LLMs into a brief statement, making the prompt and the entire input context as succinct as possible, thus guaranteeing a higher inference efficiency. Template formatting design can be divided into two parts: instruction generation and multi-step reasoning.
@@
Instruction Generation. The instruction of the template refers to a short description of the task. By adding instructions to the prompt, LLMs can quickly understand the context and the task they are currently performing, and thus may require fewer demonstrations to create a desirable prompt. The performance of a given task is highly affected by the quality of the instructions. The instructions vary not only between different datasets for the same task but also between different models. Unlike demonstrations that are usually included in traditional datasets, the generation of instructions is heavily dependent on human efforts. To enhance the efficiency of instruction generation, automatic instruction generation techniques have been proposed. For example, Instruction Induction (Honovich et al., 2022) and Automatic Prompt Engineer (Zhou et al., 2023b) have demonstrated that LLMs can generate task instructions. Wang et al. (2023b) propose Self-Instruct, an approach that allows LLMs to align with self-generated instructions, highlighting their inherent adaptability. Yang et al. (2023c) also discover that LLMs can be treated as an optimizer to iteratively generate better instructions for the target LLM and have applied this technique to various LLMs. Chen et al. (2023c) develop TeGit for training language models as task designers, which can automatically generate inputs and outputs together with high-quality instructions to better filter the noise based on a given human-written text for fine-tuning LLMs. Despite the promise of automatic instruction generation methods, their complexity is still a major bottleneck for their real-world adoption.
@@
Multi-Step Reasoning. Guiding the LLMs to produce a sequence of intermediate steps before outputting the final answer can greatly improve the quality of the generation. This technique is also referred to as Chain-of-Thought (CoT) prompting (Wei et al., 2022b). Rather than repeatedly choosing a few exemplary examples to make the context and task more understandable to the LLMs, CoT only concentrates on a limited number and adds the details for contemplation into the context, making the prompt more comprehensive and effective and guaranteeing a more efficient inference. However, despite the advantages of CoT, it is still difficult to ensure the accuracy of every intermediate step (Dong et al., 2023). Given that, many techniques have been proposed to address this issue. For example, Auto-CoT (Zhang et al., 2023f) proposes to generate the CoT step by step from LLMs. Self-Ask (Press et al., 2023) incorporates the self-generated question of each step into the CoT. ReAct (Yao et al., 2023e) performs dynamic reasoning to create, maintain, and adjust high-level plans for acting, while interacting with external environments to incorporate additional information into reasoning. Least-to-Most Prompting (Zhou et al., 2023c) breaks down the complex question into smaller ones and answers them iteratively within the context of former questions and answers. Tree-of-Thought (ToT) (Yao et al., 2023c) expends CoT to include exploration over coherent units of text and deliberates decision-making processes. CoT-SC (Wang et al., 2022e) introduces a novel decoding approach called “self-consistency” to replace the simplistic greedy decoding in CoT prompting. It starts by sampling various reasoning paths instead of just the greedy one and then determines the most consistent answer by considering all the sampled paths. Graph of Thoughts (GoT) (Besta et al., 2023) represent information produced by an LLM as a generic graph, with “LLM thoughts” as vertices and edges indicating dependencies between these vertices. Contrastive CoT (Chia et al., 2023) proposes contrastive chain of thought to enhance language model reasoning by providing both valid and invalid reasoning demonstrations. Lastly, XoT (Ding et al., 2023b) utilizes pretrained reinforcement learning and Monte Carlo Tree Search (MCTS) to integrate external domain knowledge into LLMs’ thought processes, thereby boosting their ability to efficiently generalize to new, unseen problems.
3.2.2. Prompt Compression

### v4

@@ 2.5.4 Transformer-Alternate Architectures
State Space Models. A promising approach that aims to substitute the attention mechanism is state space models (SSMs). SSM is formulated as , , which maps a single-dimension input signal to an N-dimension latent state before projecting to a single-dimension output signal , where , , , are parameters learned by gradient descent (Gu et al., 2022a). Compared to attention that has quadratic complexity, SSMs provide near-linear computational complexity related to the length of the sequence. Given such advantage, a series of techniques have been proposed to improve SSMs. For example, the Structured State Space (S4) sequence model (Gu et al., 2022a) refines SSMs by conditioning matrix with a low-rank correction. This enables stable diagonalization and simplifies the SSM to the well-studied computation of a Cauchy kernel. Diagonal State Space (DSS) (Gupta et al., 2022) improves SSMs by proposing fully diagonal parameterization of state spaces instead of a diagonal plus low rank structure. To bridge the gap between SSMs and attention while adapting to modern hardware, H3 (Hungry Hungry Hippo) (Fu et al., 2023a) stacks two SSMs to interact with their output and input projection, allowing it to log tokens and facilitate sequence-wide comparisons. Mehta et al. (2023) introduce a more efficient layer called Gated State Space (GSS), which has been empirically shown to be 2 to 3 times faster than DSS while maintaining the perplexity on multiple language modeling benchmarks. Block-State Transformer (BST) (Pilault et al., 2023) designs a hybrid layer that combines an SSM sublayers for extended range contextualization with a Block Transformer sublayer for short-term sequence representation. Gu & Dao (2023) propose Mamba to enhance SSMs by designing a selection mechanism to eliminate irrelevant data and develop a hardware-aware parallel algorithm for recurrent operation, achieving 5x throughput than Transformers. Ren et al. (2023a) extend SSMs and propose a general modular activation mechanism named Sparse Modular Activation (SMA), which unifies MoE, adaptive computation, dynamic routing and sparse attention, and further applies SMA to develop a novel architecture, SeqBoat, to achieve state-of-the-art quality-efficiency trade-off.
@@
Other Sequential Models. Some other architectures have been proposed to replace the Transformer layer. For instance, Receptance Weighted Key Value (RWKV) model (Peng et al., 2023b) combines the advantages of recurring neural networks (RNN) and Transformers. Such combination utilizes the effective parallelizable training feature of Transformers coupled with the efficient inference ability of RNNs, thereby effectively tackling the challenges associated with long sequence processing, outperforming Transformer-based models such as BLOOM and OPT. Poli et al. (2023) propose Hyena, a sub-quadratic alternative to the attention mechanism to mitigate the quadratic cost in long sequences. This operator includes two efficient sub-quadratic primitives: an implicit long convolution and multiplicative element-wise gating of the input. Hyena facilitates the development of larger, more efficient convolutional language models for long sequences and outperforms RWKV and GPT-Neo on SuperGLUE tasks (Wang et al., 2019). Lastly, MEGABYTE (YU et al., 2023) breaks down long sequences into fixed-sized patches akin to tokens, comprising a patch embedder for encoding, a global module acting as a large autoregressive Transformer for patch representations, and a local module for predicting bytes within a patch.
3 Data-Centric Methods
@@
where level=1text width=22em,font=,,
@@
Yao et al. (2022a),
DSIR (Xie et al., 2023b),
DoReMi (Xie et al., 2023a), leaf, text width=44em
@@
Ivison et al. (2023),
Instruction Mining (Cao et al., 2023),
TS-DShapley (Schoch et al., 2023),
LTD Instruction Tuning (Chen et al., 2023b),
AlpaGasus (Chen et al., 2024b),
LIMA (Zhou et al., 2023a), leaf, text width=47em
3.1 Data Selection
Data selection is a fundamental technique for enhancing efficiency. As summarized in Figure 16, in the context of LLMs, data selection techniques have been primarily used for enhancing the efficiency of pre-training and fine-tuning.
3.1.1 Data Selection for Efficient Pre-Training
@@
Data selection enhances LLMs pre-training efficiency by strategically selecting informative and diverse data samples during training. For example, SSPT (Glass et al., 2020) is a pre-training task based on the principles of reading comprehension. It involves selecting answers from contextually relevant text passages, which has shown notable improvements in performance across various Machine Reading Comprehension benchmarks. Yao et al. (2022a) propose a meta-learning-based method for selecting linguistically informative sentences which significantly elevates the quality of machine-generated translations. Xie et al. (2023b) propose DSIR, a data selection method based on importance resampling for both general-purpose and specialized LLMs. It calculates how important different pieces of data are within a simpler set of features and chooses data based on these importance calculations. Experimental results demonstrate that DSIR achieves similar performance to expert curation across eight different target distributions. In the context of pre-training general-domain models, DSIR outperforms random selection and heuristic filtering baselines by 2–2.5% on the GLUE benchmark. Different from DSIR, Xie et al. (2023a) design DoReMi to address the distribution shift between pre-training and downstream tasks, which is also a critical problem for training data selection.
3.1.2 Data Selection for Efficient Fine-Tuning
@@
Data selection can also boost LLM fine-tuning efficiency given that only a curated subset of examples is employed to refine the model. For example, Ivison et al. (2023) propose to use a few unlabeled data samples to retrieve similar labeled ones from a larger multitask dataset, improving task-specific model training. This method outperforms standard multitask data sampling for fine-tuning and enhances few-shot fine-tuning, yielding an 2-23% relative improvement. With the success of instruction tuning, many studies start focusing on the selection of high-quality instruction data to fine-tune LLMs. For example, Instruction Mining (Cao et al., 2023) presents a linear evaluation method to assess data quality in instruction-following tasks. It highlights the importance of high-quality data, showing that models trained with Instruction Mining-curated datasets outperform those trained on generic datasets in 42.5% of the considered cases. TS-DShapley (Schoch et al., 2023) is introduced to address the computational challenges of applying Shapley-based data valuation to fine-tuning LLMs. It employs an efficient sampling-based method that aggregates Shapley values computed from subsets to evaluate the entire training set. Low Training Data Instruction Tuning (LTD Instruction Tuning) (Chen et al., 2023b) challenges the need for large datasets in fine-tuning, showing that less than 0.5% of the original dataset is able to effectively train task-specific models without compromising performance. This approach enables more resource-efficient practices in data-scarce environments, combining selective data strategies with tailored training protocols for optimal data efficiency. AlpaGasus (Chen et al., 2024b) is a model fine-tuned on a mere 9K high-quality data samples, which are meticulously filtered from a larger dataset of 52K. It outperforms the original model trained on the full dataset and reduces the fine-tuning time by 5.7x, demonstrating the power of high-quality data in instruction-fine-tuning. Lastly, LIMA (Zhou et al., 2023a) fine-tunes LLMs with a small, selected set of examples, showing strong performance and challenging the need for extensive tuning. It generalizes well to new tasks, matching or exceeding GPT-4 in 43% of the considered cases.
3.2 Prompt Engineering
Prompt engineering (Liu et al., 2023a) focuses on designing effective inputs (i.e., prompts) to guide LLMs in generating desired outputs. It enhances inference efficiency by tailoring the input prompts or queries to better suit the capabilities of a specific language model. When used for some simple tasks, such as semantic classification, prompt engineering can even substitute fine-tuning to achieve high accuracy (Liu et al., 2022a). As summarized in Figure 18, prompt engineering techniques can in general be grouped into few-shot prompting, prompt compression, and prompt generation.
3.2.1 Few-Shot Prompting
Few-shot prompting aims to provide a LLM with a limited set of examples, referred to as demonstrations, to steer its understanding to a task it is required to execute (Wei et al., 2022a). These demonstrations are selected from the training corpus based on their similarity to the test example, and the LLM is expected to use the knowledge gained from these similar demonstrations to make the correct prediction (Dong et al., 2023). Few-shot prompting provides an efficient mechanism to use LLM by guiding the LLM to perform a wide variety of tasks without the need for additional training or fine-tuning. Furthermore, an effective few-shot prompting approach can make the created prompt concise enough to allow LLMs to quickly adjust to the task in high accuracy with only a slight increase of extra context, thus significantly improving inference efficiency. As illustrated in Figure 19, few-shot prompting techniques can in general be grouped into demonstration organization and template formatting.
@@ 3.2.1 Few-Shot Prompting
where level=1text width=12em,font=,,
where level=2text width=15em,font=,,
where level=3text width=13em,font=,,
@@
KATE (Liu et al., 2022b),
VoteK (SU et al., 2023),
Wang et al. (2023f),
IDS (Qin et al., 2023a),
Min et al. (2022b),
LENS (Li & Qiu, 2023),
MDL (Wu et al., 2023c),
Zhang et al. (2022b),
EPR (Rubin et al., 2022),
UDR (Li et al., 2023f),
Wang et al. (2024b),
@@
, leaf, text width=32em
@@
Lu et al. (2022),leaf, text width=7em
@@
Instruction Induction (Honovich et al., 2023),
Automatic Prompt Engineer (Zhou et al., 2023c),
Self-Instruct (Wang et al., 2023h),
OPRO (Yang et al., 2023a),
TeGit (Chen et al., 2023h)
, leaf, text width=37em
@@
ReAct (Yao et al., 2023b),
Least-to-Most Prompting (Zhou et al., 2023b),
Tree-of-Thought (Yao et al., 2023a),
CoT-SC (Wang et al., 2023h),
Graph of Thoughts (Besta et al., 2024),
Contrastive CoT (Chia et al., 2023),
XoT (Ding et al., 2023b),
Skeleton-of-Thought (Ning et al., 2024), leaf, text width=44em
@@
Gisting (Mu et al., 2023), AutoCompressors (Chevalier et al., 2023),
PCRL (Jung & Kim, 2023),
ICAE (Ge et al., 2024),
Nugget 2D (Qin et al., 2023b),
LongLLMLingua (Jiang et al., 2024), leaf, text width=43em
@@
TempLM (Zhang et al., 2023e),
PromptGen (Zhang et al., 2022c),
leaf, text width=44em
@@
Demonstration Organization. Demonstration organization refers to organizing the demonstrations in an appropriate way so as to form a suitable prompt for inference. Demonstration organization has a significant impact on the inference efficiency since improper organization may result in the processing of a considerable amount of unnecessary information, leading to significant slowdown. The optimization of demonstration organization comes from its two main steps: demonstration selection and demonstration ordering.
@@
Demonstration Selection. Demonstration selection aims to choose the good examples for few-shot prompting (Dong et al., 2023). In order to generate a satisfactory result, a good selection of demonstrations may only require a few number of demonstrations to be used for the prompt, thus making the prompt concise and straightforward for a more efficient inference. Existing demonstration selection techniques can be grouped into unsupervised methods (Liu et al., 2022b; SU et al., 2023; Wang et al., 2023f; Qin et al., 2023a; Min et al., 2022b; Li & Qiu, 2023; Wu et al., 2023c; Zhang et al., 2022b) and supervised methods (Rubin et al., 2022; Li et al., 2023f; Wang et al., 2024b; Luo et al., 2023). Unsupervised methods aim to select the nearest examples from the training set using a predefined similarity function, such as L2 distance, cosine distance, and minimum description length (MDL) (Wu et al., 2023c). For example, KATE (Liu et al., 2022b) is an unsupervised selection method that directly uses the nearest neighbors of a given test sample as the corresponding demonstrations. VoteK (SU et al., 2023) is an improved version of KATE to resolve its limitation that requires a large set of examples to achieve good performance. Unlike KATE, VoteK increases the diversity of the demonstrations by penalizing examples similar to those already selected. In contrast, supervised methods require training a domain-specific retriever from the training set and using it for demonstration selection. For example, EPR (Rubin et al., 2022) is trained to select demonstrations from a small set of candidates initialized by the unsupervised retriever such as BM25 from the training corpus. UDR (Li et al., 2023f) further enhances EPR by adopting a unified demonstration retriever to unify the demonstration selection across different tasks. Compared to unsupervised methods, supervised methods often lead to a more satisfying generation result but require frequent adjustment of the retriever for handling the out-of-domain data, making them relatively less efficient for inference.
@@
Demonstration Ordering. After selecting representative samples from the training set, the next step is ordering these samples in the prompt. The order of the demonstrations also has a significant impact on the performance of the model. Therefore, selecting the right order of demonstrations can help the model quickly reach a good generation quality with fewer samples, thus improving the inference efficiency. To date, only a few studies have delved into this area. For example, Liu et al. (2022b) suggest arranging demonstrations based on their distance from the input, placing the closest demonstration furthest to the right. Lu et al. (2022) propose to develop both global and local entropy metrics and use the entropy metrics to set up the demonstration order.
@@
Template Formatting. Template formatting aims to design a suitable template to form the prompt. A good template typically compiles all the information needed by LLMs into a brief statement, making the prompt and the entire input context as succinct as possible, thus leading to a higher inference efficiency. Template formatting can be divided into two parts: instruction generation and multi-step reasoning.
@@
Instruction Generation. The instruction of the template refers to a short description of the task. By adding instructions to the prompt, LLMs can quickly understand the context and the task they are currently performing, and thus may require fewer demonstrations to create a desirable prompt. The performance of a given task is highly affected by the quality of the instructions. The instructions vary not only between different datasets for the same task but also between different models. Unlike demonstrations that are usually included in traditional datasets, the generation of instructions is heavily dependent on human efforts. To enhance the efficiency of instruction generation, automatic instruction generation techniques have been proposed. For example, Instruction Induction (Honovich et al., 2023) and Automatic Prompt Engineer (Zhou et al., 2023c) demonstrate that LLMs can generate task instructions. Wang et al. (2023h) propose Self-Instruct, an approach that allows LLMs to align with self-generated instructions, highlighting their inherent adaptability. Experimental results show that it achieves an 33% improvement over the original model when applied on vanilla GPT3. Yang et al. (2023a) also discover that LLMs can be treated as an optimizer to iteratively generate better instructions for the target LLM and have applied this technique to various LLMs. Experiments demonstrate that the best prompts optimized by this method outperform human-designed prompts by up to 8% on GSM8K, and by up to 50% on Big-Bench Hard tasks. Chen et al. (2023h) develop TeGit for training language models as task designers, which can automatically generate inputs and outputs together with high-quality instructions to better filter the noise based on a given human-written text for fine-tuning LLMs. Despite the promise of automatic instruction generation methods, their complexity is still a major bottleneck for their real-world adoption.
@@
Multi-Step Reasoning. Multi-step reasoning (Huang & Chang, 2023) refers to techniques that guide the LLMs to produce a sequence of intermediate steps before outputting the final answer. Compared to fine-tuning, conducting specific task reasoning directly through this way is a more efficient approach. At the same time, its accuracy improvement is often not as good as fine-tuning. One of the widely used techniques in multi-step reasoning is Chain-of-Thought (CoT) prompting (Wei et al., 2022b), which adds a series of reasoning steps into the prompt to make it more informative and comprehensive. Despite its advantages, it is still difficult for CoT to ensure the accuracy of every intermediate step (Dong et al., 2023). A number of techniques have been proposed to address this issue. For example, Auto-CoT (Zhang et al., 2023g) proposes to generate the CoT step by step from LLMs. Self-Ask (Press et al., 2023) incorporates the self-generated questions of each step into CoT. ReAct (Yao et al., 2023b) performs dynamic reasoning to create, maintain, and adjust high-level plans for acting, while interacting with external environments to incorporate additional information into reasoning. Least-to-Most Prompting (Zhou et al., 2023b) is a new milestone in CoT. It breaks down the complex question into smaller ones and answers them iteratively within the context of former questions and answers. Experimental results show that Least-to-Most Prompting can boost the accuracy of GPT-3 code-davinci-002 model with CoT prompting to 99.7% by using just 14 exemplars. Tree-of-Thought (ToT) (Yao et al., 2023a) expends CoT to include exploration over coherent units of text and deliberates decision-making processes. It outperforms GPT-4 with CoT prompting in the Game of 24 benchmark. CoT-SC (Wang et al., 2023h) introduces a novel decoding approach called self-consistency to replace the greedy decoding in CoT prompting. It starts by sampling various reasoning paths instead of just the greedy one and then determines the most consistent answer by considering all the sampled paths. Graph of Thoughts (GoT) (Besta et al., 2024) represents information produced by an LLM as a generic graph, with “LLM thoughts” as vertices and edges indicating dependencies between these vertices. Experimental results show that GoT increases the quality of sorting by 62% over ToT while reducing the cost by over 31%. Contrastive CoT (Chia et al., 2023) proposes to enhance language model reasoning by providing both valid and invalid reasoning demonstrations. XoT (Ding et al., 2023b) utilizes pretrained reinforcement learning and Monte Carlo Tree Search (MCTS) to integrate external domain knowledge into the thought processes of LLMs, thereby boosting their ability to efficiently generalize to new, unseen problems. Lastly, Skeleton of Thought (SoT) (Ning et al., 2024) proposes a method that first prompts the LLM to organize the output and then parallelizes the generation of different segments. Through splitting the serial generation into two different steps, it makes the generation workload more parallelizable. Therefore, it improves hardware utilization and provides speedups across LLM, revealing the possibility of exploiting data-level organization to enhance efficiency.
3.2.2 Prompt Compression